6 defects are observed in 283 unites produced, I need to determine the probability of obtaining units with zero defects (yield)

6/283=.021201
.021201
this is where i am stuck. is this binomial ?

Answer by stanbon(75887)

6 defects are observed in 283 unites produced, I need to determine the probability of obtaining units with zero defects (yield)
6/283=.021201
.021201
==================
It's binomial because each unit is either defective or not defective.
-----
6/283 is the probability of a unit being defective.
----
1 - (6/283) is the probability of a unit not being defective.
----
1 - (6/283) = 1 - 0.021201 = 0.9788
==================================
Cheers,
Stan H.
=============